Finding 3 consecutive positive integers such the product of the first and the third is 60.
Answer by Alan3354(69443)

Finding 3 consecutive positive integers such the product of the first and the third is 60.
---------------
Those would be n and n+2.
n*(n+4) = 60
n^2 + 4n - 60 = 0
(n+10)*(n-6) = 0
n = 6